# A time-to-digital converter with steady calibration through single-photon detection

### v1

@@ II.3 Temperature test setup
The validity of the steady calibration method was verified by forcing a change in the working point of the system, which should require a calibration. Therefore, the device was tested at different temperatures as this can have a relevant impact on the performance of a TDC and is considered strictly related to the TDC calibration process [24]. To modify the device temperature, we used a combination of a Peltier cell and an RAL 9006 climatic chamber by Angelatoni Test Technologies. The combination of both temperature control systems allowed the FPGA chip to reach a broader temperature range, from 5 °C to 80 °C. To measure the temperature of the chip, the included XADC sensor of the board was used, which gives a temperature measurement in the center of the chip itself.
The jitter and resolution of a two-channel implementation of MARTY was used to characterize its performance. To calculate the jitter, an input signal is divided into the two channels, which we assume equal for this purpose, and the histogram of the differences between the timetags is obtained, from which the jitter can be obtained. For this approximation, we assume that both channels have the same jitter and that the jitter of the source does not affect the resulting histogram due to both channels receiving the same input signal.
To perform the temperature measurements, a 1550 nm attenuated laser source was used as an input to a single-photon detector (PMD-IR detector by MPD), which was then divided and used as input for both channels of the TDC. The chip temperature was scanned in the whole range of 5 °C to 80 °C with a step of 1 °C, where for each temperature the jitter and resolution were evaluated.
@@ II.4 Quantum Key Distribution test setup
To evaluate the device performance for QKD applications, while also comparing it with commercial TDCs, a QKD transmitter like the one proposed in [25] was used to encode an “HVDD” sequence. [^3] An “HVDD” sequence is a sequence of four qubits encoded using single-photon with Horizontal-Vertical-Diagonal-Diagonal polarizations. We used a standard polarization receiver scheme with four PMD-IR detectors by MPD, where for the time-tagger we used both a QuTAG from QuTools and MARTY. To evaluate the performance, a detection peak was measured with both devices, and the Quantum Bit Error Rate (QBER) for the HVDD sequence was estimated.
@@ III Results and discussions
Figure 3: DNL (in blue) and $\text{DNL}^{t}$ (in red) obtained for a single channel of the TDC at room temperature.
@@
Figure 4: Calibrated time as a function of the bin number of the TDC, in orange the ideal behaviour for a uniform delay line.
@@
MARTY’s main performance metrics, such as resolution, differential nonlinearity (DNL), and jitter were characterized using a two-channel TDC configuration. The DNL represents the deviation of the delay line with respect to an ideal one (uniformly distributed in time) and it was evaluated through a code density test with an internal RO as the input of the TDC, obtaining the counts for bin $i$, $w_{i}$ (Fig. 2a), from which the DNL was obtained as
$\text{DNL}_{i}=\frac{w_{i}-\langle w\rangle}{\langle w\rangle},$ (3)
[…]
$\tau_{\mathrm{res}}=\frac{t^{(c)}_{N_{c}}}{N_{c}}=18.22~{}\text{ps}$ (4)
was obtained at room temperature. Moreover, once the bins were properly calibrated, the DNL was evaluated again, this time using calibrated times $t^{(c)}$ as
${\rm DNL}^{t}_{i}=\begin{cases}0&\quad\text{if i=1},\\ \frac{t^{(c)}_{i}-t^{(c)}_{i-1}-\langle\Delta t^{(c)}\rangle}{\langle\Delta t^{(c)}\rangle}&\quad\text{if i}>1,\\ \end{cases}$ (5)
where $\langle\Delta t^{(c)}\rangle=\langle t_{i+1}^{(c)}-t_{i}^{(c)}\rangle=\tau_{\mathrm{res}}$.
@@
Lastly, to characterize the jitter of the system, a scheme like the one described in section II.3 was used, paired with a square wave as an input for both channels at room temperature. To obtain the jitter, a Gaussian fit was applied to the histogram of the difference between consecutive tags, from which the jitter was estimated as the $\text{FWHM}=27.11$ ps (Fig. 5). We notice that the jitter is obtained by dividing by $\sqrt{2}$ the standard deviation of the Gaussian fit.
@@
Figure 5: Histogram of the difference between two consecutive timetags in a two-channel implementation. In red, the corresponding Gaussian fit, from which the jitter was estimated.
@@
To determine the device performance with respect to environment changes, a temperature characterization of the system using the setup described in Section II.3 was performed. The data acquisition process consisted on storing, for each temperature, a file with the calibration curve for each channel along with the registered tags in non-calibrated format, with their respective channel. This format allowed to analyze the data with or without calibration, while also letting the possibility to implement calibration at particular points in the measurement. The acquired data for each temperature was processed in four different ways:
@@
    Calibrated using the calibration curve obtained with the RO at 5 °C (light orange curve in Fig. 7).
@@
    Calibrated using the first time tags registered with the single photon detector at 5 °C (light blue curve in Fig. 7).
@@
Figure 7: Comparison of the jitter of the system as a function of temperature obtained for different calibration methods: with the RO at 5 °C and maintained through the scan (light orange); with the RO every 1 °C step (red); with the single photon detector at 5 °C and maintained through the scan (light blue); and steadily with the single photon detector (blue).
@@
As a first step, the relation between temperature and the average propagation time inside the delay line, $\langle\delta t\rangle$, was analysed. Since the maximum number of registered bins is temperature and channel dependent (obtaining around 130 bins for 5 °C and up to 136 for 80 °C), we restricted this analysis to the first 120 elements of the delay line, such that
$\langle\delta t\rangle=\frac{\sum_{j=1}^{120}\delta t_{j}}{120}.$ (6)
@@
The observed behavior depends on the particular architecture and materials of the Zynq7020 chip (28 nm process), of which not much information is publicly available, concluding that the CARRY4 element has a propagation time that is, in average, inversely proportional to temperature.
@@
Then, the jitter was calculated in the 4 scenarios previously mentioned (Fig. 7). For both cases in which the calibration was done just at 5 °C, increasing the chip temperature monotonically increased the jitter of the system (as also verified by Alshehry et al. in [26]), which highlights the importance of the steady calibration scheme, since the transistors’ characteristics are temperature dependent. On the other hand, when applying a calibration for every temperature step, the jitter decreased until 30 °C, which implies an improvement in performance, and then increased until it stabilized at 60 °C, close to the starting value. Finally, the one obtained with the steady calibration had a very similar shape to the one obtained by calibrating at each temperature, but with two notable improvements. In first instance, the system’s jitter was found to be more stable, presenting an average standard deviation $\langle\sigma_{\text{steady}}\rangle=0.64$ ps, compared to the $\langle\sigma_{\text{RO}}\rangle=1.33$ ps obtained for the case when RO calibration was performed before each measurement. Moreover, as previously mentioned, the steady calibration does not require the acquisition to be stopped at any time, thus allowing the system to not lose any data.
@@
Finally, a test with a QKD polarization-encoded source and receiver scheme was performed. The width of the detected pulse depends on the width of the optical pulse itself, the jitter of the detector, and the jitter of the time tagger. For both devices, the mean QBER obtained was of $2.2$ [^4] The QKD source and receivers were not properly optimized for performance as it is outside the scope of this work., showing that our design presents good performance for QKD applications. We compared the width of the detected pulse using both MARTY and the QuTAG, a high-level commercial time tagging system with 13.4 ps jitter on the standard version (Fig. 8). It was observed that, even though Marty’s jitter was higher than the QuTAG, the pulse width obtained with both devices were comparable. This behaviour is due to most of the width of the pulse coming from the detector’s jitter, which is considerable higher than the time-tagger ($\sim 100$ ps). Indeed, moving to a higher level FPGA technology (e.g., Ultrascale+ with 16 nm FinFET process) can considerably increase the device’s performances [27].

### v2

@@ II.3 Temperature test setup
The validity of the steady calibration method was verified by forcing a change in the working point of the system, which should require a calibration. Therefore, the device was tested at different temperatures as this can have a relevant impact on the performance of a TDC and is considered strictly related to the TDC calibration process [7]. By modifying the temperature of the system in a controlled and fast manner, the reliability of the calibration method against changes in the environment conditions was tested. This is equivalent to testing the system over extended periods of time on a slower-changing environment. To modify the device temperature, we used a combination of a Peltier cell and an RAL 9006 climatic chamber by Angelatoni Test Technologies. The combination of both temperature control systems allowed the FPGA chip to reach a wider temperature range, from 5 C to 80 C  [^5] The scanned range could not be further increased due to the climatic chamber not including humidity control. Thus, to protect the board integrity, we remained over 0 C. To measure the temperature of the chip, the included XADC sensor of the board was used, which gives a temperature measurement in the center of the chip itself.
The FWHM jitter and resolution of a two-channel implementation of MARTY was used to characterize its performance. To calculate the FWHM jitter, an input signal is divided into the two channels, which we assume equal for this purpose, and the histogram of the differences between the timetags is obtained, from which the FWHM jitter can be obtained. For this approximation, we assume that both channels have the same FWHM jitter and that the jitter of the source does not affect the resulting histogram due to both channels receiving the same input signal.
To perform the temperature measurements, a 1550 nm attenuated laser source was used as an input to a single-photon detector (PMD-IR detector by MPD), which was then divided and used as input for both channels of the TDC. The chip temperature was scanned in the whole range of 5 C to 80 C with a step of 1 C, where for each temperature the FWHM jitter and resolution were evaluated.
@@ II.4 Quantum Key Distribution test setup
To evaluate the device performance for QKD applications, while also comparing it with commercial TDCs, a QKD transmitter using a three-state one-decoy BB84 protocol at $R=50$ MHz repetition rate, based on the source presented in [5], was used to encode an “HVDD” sequence [^6] An “HVDD” sequence is a sequence of four qubits encoded using single-photon with Horizontal-Vertical-Diagonal-Diagonal polarizations.. We used a standard polarization receiver scheme with four PMD-IR detectors by MPD, where for the time-tagger we used both a QuTAG from QuTools and MARTY. To evaluate the performance, a detection peak was measured with both devices, and the Quantum Bit Error Rate (QBER) for the HVDD sequence was estimated.
@@ III Results and discussions
Figure 4: DNL (in orange) and tDNL (in blue) obtained for a single channel of the TDC at room temperature.
@@
Figure 5: Calibrated time $t^{(c)}$ as a function of the bin number of the TDC, in orange the ideal behaviour for a uniform delay line.
MARTY’s main performance metrics, such as resolution, differential nonlinearity (DNL), and FWHM jitter, were characterized using a two-channel TDC configuration [19]. As first instance, the mean resolution of the system was evaluated at room temperature such that
$\tau_{\mathrm{res}}=\frac{\tau_{\rm coarse}}{N_{c}}=18.37~\text{ps}.$ (5)
@@
The DNL represents the deviation of the delay line with respect to an ideal one (uniformly distributed in time) [15]. The DNL can be evaluated through a code density test with an internal RO as the input of the TDC, representing a signal uniformly distributed over the time at the input of the TDC. From the obtained counts for bin $i$, $w_{i}$ (Fig. 3a), the DNL can be estimated as
$\text{DNL}_{i}=\frac{w_{i}-\langle w\rangle}{\langle w\rangle}=\frac{\delta t_{i}-\langle\delta t\rangle}{\langle\delta t\rangle},$ (6)
where $\langle w\rangle=\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}w_{i}$, $\langle\delta t\rangle=\frac{1}{N_{c}}\sum_{i}\delta t_{i}=\tau_{\rm res}$ and $N_{c}$ is the last bin for which counts were recorded (Fig. 4). A $\text{DNL}=[-0.97,\,2.86]$ was obtained, but at room temperature only 21 out of the 132 bins had $\text{DNL}>1$. Moreover, after obtaining the calibrated times $t^{(c)}$ (Fig. 5), a non-linearity quantity, named $\rm tDNL$, is defined as
${\rm tDNL}_{i}=\frac{\delta t^{(c)}_{i}-\langle\delta t^{(c)}\rangle}{\langle\delta t^{(c)}\rangle}$ (7)
where $\delta t^{(c)}_{i}$ corresponds to the difference between consecutive calibrated time bins and $\langle\delta t^{(c)}\rangle=\tau_{\rm res}$ as defined in Section II.2. It is worth noting that ${\rm tDNL}_{i}$ does not provide additional information regarding the linearity of the system [^7] Please note that this quantity is not a DNL by the standard definition., but instead provides a linearity notion ‘as seen by the user’, i.e. a quantity that is relevant for the final user of the device, by utilizing the center of the bins instead of the edges of the bins for the DNL calculation.
@@
Lastly, to characterize the FWHM jitter of the system, a scheme like the one described in Section II.3 was used, paired with a square wave as an input for both channels at room temperature. To obtain the FWHM jitter, a Gaussian fit was applied to the histogram of the difference between consecutive tags, from which the FWHM jitter was estimated as the $\text{FWHM}=27.63$ ps (Fig. 6).
@@
Figure 6: Histogram of the difference between two consecutive timetags in a two-channel implementation. In red, the corresponding Gaussian fit, from which the FWHM jitter was estimated.
@@
To determine the performance of the device with respect to changes in the environment, a temperature characterization of the system was performed using the setup described in Section II.3. The data acquisition process consisted on storing, for each temperature, a file with the calibration curve for each channel along with the registered tags in noncalibrated format with their respective channel. This format allowed to analyze the data with or without calibration, while also letting the possibility to implement calibration at particular points in the measurement. The acquired data for each temperature was processed in four different ways:
@@
    Calibrated using the calibration curve obtained with the RO at 5 C (light orange curve in Fig. 7).
@@
    Calibrated using the first time tags registered with the single photon detector at 5 C (light blue curve in Fig. 7).
@@
Figure 7: Comparison of the FWHM jitter of the system as a function of temperature obtained for different calibration methods: with the RO at 5 C and maintained through the scan (light orange); with the RO every 1 C step (red); with the single photon detector at 5 C and maintained through the scan (light blue); and steadily with the single photon detector (blue).
@@
As a first step, the relation between temperature and the average propagation time within the delay line, $\langle\delta t\rangle$, was analyzed. Since the maximum number of registered bins is temperature and channel dependent (obtaining around 129 bins for 5 C and up to 135 for 80 C), we restricted this analysis to the first 120 elements of the delay line, such that
$\langle\delta t\rangle_{\rm truncated}=\frac{\sum_{j=1}^{120}\delta t_{j}}{120}.$ (8)
@@
The observed behavior depends on the particular architecture and materials of the Zynq7020 chip (28 nm process), of which not much information is publicly available, concluding that the CARRY4 element has a propagation time that is, on average, inversely proportional to temperature (see Fig. 8).
@@
Then, the FWHM jitter was calculated in the 4 scenarios previously mentioned (Fig. 7). For both cases in which calibration was done just at 5 C, increasing the chip temperature monotonically increased the FWHM jitter of the system (as also verified by Alshehry et al. in [2]), which highlights the importance of the steady calibration scheme, since transistors’ characteristics are temperature dependent. On the other hand, when applying a calibration for every temperature step, the FWHM jitter decreased until 30 C, which implies an improvement in performance, and then increased until it stabilized at 60 C, close to the starting value. Finally, the one obtained with the steady calibration had a very similar shape to the one obtained by calibrating at each temperature, but with two notable improvements. In first instance, the system’s FWHM jitter was found to be more stable, presenting an average standard deviation of $\langle\sigma_{\text{steady}}\rangle=0.61$  ps, compared to the $\langle\sigma_{\text{RO}}\rangle=1.33$ ps obtained for the case when RO calibration was performed before each measurement. Moreover, as previously mentioned, the steady calibration does not require the acquisition to be stopped at any time, thus allowing the system to not lose any data.
Finally, a test with a QKD polarization-encoded source and receiver scheme was performed. The width of the detected pulse depends on the width of the optical pulse itself, the jitter of the detector, and the jitter of the time tagger. In real-world applications of QKD protocols, it is a common practice to apply time-filtering over the detected quantum signals such that the background noise is reduced. Therefore, the jitter plays a relevant in order to lower the background counts, that are typically uniformly distributed in time. Indeed, the receiver jitter (that combines optical pulse width, single photon detector jitter and TDC jitter) determines the width of the temporal filter at the receiver. A large receiver jitter corresponds to a large time-selection window, increasing the noise and the QBER. For both devices, the mean QBER obtained was of $2.2$ % [^8] The QKD source and receivers were not properly optimized for performance as it is outside the scope of this work., showing that our design presents good performance for QKD applications. Even though the device was tested for a specific protocol (three-state one-decoy BB84), it is also applicable to all those QKD protocols that rely on time-tagging single photon events, such as discrete-variable including measurement device independent and twin-field protocols. The same can be said for the steady calibration, since all compatible protocols include a single-photon or weak coherent pulse source, which satisfy the signal requirement to be used for calibration purposes.
@@
We compared the width of the detected pulse using both MARTY and the QuTAG, a high-level commercial time tagging system with 13.4 ps FWHM jitter in the standard version (Fig. 9). It was observed that, even though Marty’s FWHM jitter was higher than the QuTAG, the pulse widths obtained with both devices were comparable. This behavior is due to most of the width of the pulse coming from the detector’s FWHM jitter, which is considerably higher than both time-taggers ($\sim 100$ ps). Moving this design to a higher-level FPGA technology (e.g., Ultrascale+ with 16 nm FinFET process) can considerably increase the device’s performances, which would not be beneficial when using InGaAs detectors such as the ones used in this work, but would make a noticeable difference when using superconductive nanowire single photon detectors, where the FWHM jitter is on the order of tens of picoseconds [9, 35].